OpenStudy (anonymous):

Help please Express (2-√3)^2 in the form b+c√3 where b and c are intergers

OpenStudy (anonymous):

Write this in a form which we can multiply: \[(2-\sqrt{3})(2-\sqrt{3})\] Foil this out to get: \[4-4\sqrt{3}+3\] Combine like terms: \[7-4\sqrt{3}\]

OpenStudy (anonymous):

u r a life saver man thanx a lot